Grade 11Mechanics

Two forces F1 =2i-5j-6k and F2 =-i+2j-I are acting on a body at the points (1,1,0)and(0,1,2)respectively .Find torque acting on the body about point (-1,0,1)

Sreenija Dowpati
6 Years agoGrade 11

1 Answer

Vaishnavi jayale
6 Years ago
Let r1=i+j
r2=j+k
r3=-i+k
So ,r2-r1=r3
So, f2-f1=f3
Torque is r cross f
So ans.  Is =-7i+2j+7k
Soanswer
And now torque is